OpenStudy (anonymous):

Thomas is keeping records at the city fishing tournament. The table shows the information on the number of each type of fish caught in the tournament so far. a. Estimate the probability that the next fish caught will be a perch. Show your work. b. Estimate the probability that the next fish will be either a bass or a walleye. Show your work. Fish Perch Sunfish Bass Walleye Number of Fish 32 21 14 12

OpenStudy (anonymous):

For the first part, the answer is probably 32 / (32 + 21 + 14 + 12) = 32 / 79 = .4050 Ie, the no. of Perch caught against the total number of fish caught.

OpenStudy (anonymous):

can you just give me a simple answer because that is pretty complicated

OpenStudy (anonymous):

It is as simple as it can be. Okay, its just the ratio between the no. of perch caught and the total number of fish caught. So, it's 32 / 79 and that equals 0.405063. (Probably 0.405 will be enough)

OpenStudy (anonymous):

thank you